A weather ballon is filled with Helium at sea level on a hot summer day. The volume of the ballon is 67.0 L and the temperature is 32° C and since it is at sea level the pressure is 1 atm. Show how to calculate the volume of the balloon after it rises to 30,000 ft where the temperature is -23° C and the pressure is 0.56 atm.


1
Expert's answer
2021-01-15T06:47:24-0500

We will use the combined gas law equation;

"\dfrac{P1V1}{T1}=\dfrac{P2V2}{T2}"


rearranging for V2 gives;

"V2=\dfrac{P1V1T2}{T1P2}"

P1 = 1atm

P2=0.56atm

T1=(32+273)=305K

T2=(-23+273)=250K

V1=67.0L

V2=?


"V2=\dfrac{1atmx67.0Lx250K}{305Kx0.56atm} =98.068L"